# Privacy-Preserving Personalized Federated Prompt Learning for Multimodal Large Language Models

### v2

@@ Abstract
Multimodal Large Language Models (LLMs) are pivotal in revolutionizing customer support and operations by integrating multiple modalities such as text, images, and audio. Federated Prompt Learning (FPL) is a recently proposed approach that combines pre-trained multimodal LLMs such as vision-language models with federated learning to create personalized, privacy-preserving AI systems. However, balancing the competing goals of personalization, generalization, and privacy remains a significant challenge. Over-personalization can lead to overfitting, reducing generalizability, while stringent privacy measures, such as differential privacy, can hinder both personalization and generalization. In this paper, we propose a Differentially Private Federated Prompt Learning (DP-FPL) approach to tackle this challenge by leveraging a low-rank adaptation scheme to capture generalization while maintaining a residual term that preserves expressiveness for personalization. To ensure privacy, we introduce a novel method where we apply local differential privacy to the two low-rank components of the local prompt, and global differential privacy to the global prompt. Our approach mitigates the impact of privacy noise on the model performance while balancing the tradeoff between personalization and generalization. Extensive experiments demonstrate the effectiveness of our approach over other benchmarks.
@@ 1 Introduction
To tackle the above question, we proposed a Differentially Private Federated Prompt Learning (DP-FPL) approach that leverages Low-Rank Adaption (LoRA) scheme and DP as part of the prompt learning process. In our framework, each client simultaneously learns a global prompt and a local prompt. The global prompt is shared in a FL manner for generalized knowledge transfer, while the local prompt is retained at each client site for personalization. Our contributions are threefold.
  • We propose a privacy-preserving personalized federated prompt learning approach with Differential Privacy for multimodal LLMs. We adopt LoRA to decompose the local prompt into two lower rank components with an additional residual term. The factorized low-rank components allow the model to capture broader patterns that are beneficial across different data distributions, aiding the generalization capability of each client. The residual term is crucial for retaining the expressiveness lost during the decomposition process, thereby preserving the client-specific learning and improving personalization.
@@ 2 Related Work
Personalized Federated Learning. There are several existing approaches that aim to learn personalized models for clients in FL settings, including clustering (Ghosh et al., 2020; Berlo et al., 2020; Shahid et al., 2021), regularization (Shoham et al., 2019; Dinh et al., 2020; Li et al., 2020) and knowledge distillation (Li & Wang, 2019; He et al., 2020; Fang & Ye, 2022). Personalized FL is most commonly approached as a multi-task learning problem that simultaneously learns two models for each client: a global model for generalized knowledge and a local model for personalized data. Existing methods accomplish this by decoupling the model parameters or layers into global and local learning components (Arivazhagan et al., 2019; Deng et al., 2020; Zhang et al., 2020; Collins et al., 2021; Jeong & Hwang, 2022; Zhang et al., 2023). In the existing literature on personalized FL, private multi-task learning approaches aim to protect training data by retaining personalized parameters and sharing differentially private generalized parameters. Examples include Jain et al. (2021), Hu et al. (2021b), Bietti et al. (2022), Yang et al. (2023b) and Xu et al. (2024). However, these methods are designed for full model training and cannot be directly applied to prompt tuning due to the difference in the parameter space. Sun et al. (2024) proposed a similar approach to ours where they incorporate LoRA with DP in a standard FL setting, but they do not consider personalization and prompt learning, making their method not applicable to our setting.
@@ 3 Proposed Method
We introduce our proposed method, Differentially Private Federated Prompt Learning (DP-FPL), shown in Figure 1. Our approach leverages LoRA with an additional residual term to balance personalization and generalization in a differentially private FPL system.
@@ 3.1 Preliminaries on Personalized Federated Prompt Learning
An example of the personalized prompt is illustrated in Figure 1, in which each client has a collection of images captured from different angles, reflecting the heterogeneity in their local data. Consequently, client $1$ might use the prompt ”an upside-down photo of [class]”, while client $i$ could have the prompt ”a upright photo of [class]”, and client $N$ might use ”a rotated photo of [class]”. In this instance, the template ”a photo of [class]” is the generalized global prompt shared across clients, and the terms ”upside-down”, ”upright” and ”rotated” represent the personalized characteristics of each client’s prompt. These variations in prompts allow the client model to better adapt to the specific characteristics of their data, improving personalization.
@@ 3.2 Balancing Personalization and Generalization
Prior research (Aghajanyan et al., 2020; Cui et al., 2024) has demonstrated that fine tuning pre-trained LLMs with lower dimension reparameterization promotes generalization capability across various tasks. Additionally, lower intrinsic dimension improves the model utility under the effect of DP noise (Yu et al., 2021; Xu et al., 2024). Inspired by this, we adopt Low-Rank Adaptation (LoRA) (Hu et al., 2021a) in our FPL framework to balance the personalization and generalization learning under the influence of DP noise.
LoRA has been used to fine-tune pre-trained models more efficiently by introducing rank decomposition into the weight matrices of neural networks. However, LoRA has difficulty matching the performance of full-rank training in many difficult tasks (Liu et al., 2024; Biderman et al., 2024; Ivison et al., 2023; Zhuo et al., 2024). This is because low-rank decomposition methods restrict the parameter space, throwing away some of the information of the full-rank space (Konečnỳ, 2016). Recent work Cui et al. (2024) also used the low-rank approximation in non-private FPL systems, however, they factorize the local prompt only once at the beginning and train iteratively with the low-rank components. This approach can reduce the overall expressive power over the training phase. As we will show in Section 4, the loss of expressiveness can amplify the adverse effects of the DP noise, further diminishing the model’s performance. To overcome this issue, we perform the factorization process in every training round rather than only at the beginning like Cui et al. (2024), and we incorporate a residual term to compensate for the lost expressiveness. We analyze in detail the benefit of the residual term in Section 3.3. The overall personalized prompt of client $i$ can then be expressed as $p_{i}=p_{G,i}+u_{i}v_{i}+r_{i}$, where $u_{i}$ and $v_{i}$ are the factorized low-rank components and $r_{i}$ is the additional residual term.
We utilize a version of the Reparametrized Gradient Perturbation (RGP) method introduced in Yu et al. (2021) as our LoRA decomposition scheme. We perform the factorization process for each client local prompt $p_{L,i}$ in every training iteration to get the temporary low-rank prompt components $u_{i}$ and $v_{i}$. We also compute a residual term $r_{i}$ which is the remainder of the factorization process, i.e. $r_{i}=p_{L,i}-u_{i}v_{i}$. As shown by Yu et al. (2021), each client can reconstruct the gradient of $p_{L,i}$ using the gradient of the low-rank terms in back propagation as follows:
@@ 3.3 Preserving Privacy
The two main building blocks in our approach, i.e., low-rank decomposition and DP, naturally introduce error to the training process. We conjecture that this error acts as a regularization term that prevents clients from overfitting to local data, reducing personalization and improving generalization. However, under strictly private conditions (lower rank and higher DP noise), the accumulated error may become too large and potentially destroy the personalization capability. In this case, the added residual term compensates for the regularization-like error and helps improve local learning, balancing personalization and generalization. We demonstrate the benefit of the residual term in the ablation study in Section 4.3, supporting our hypothesis. Further theoretical analysis of the residual term can be a potential future work direction.
@@ 3.4 Algorithm
Low-rank decomposition via traditional SVD method requires significant runtime. Instead, we use the power method with one iteration, significantly reducing the computational cost. Given the full-rank matrix of size $m\times n$ (assuming $m\leq n$), the computational cost of SVD scales with $\mathcal{O}(m^{2}n)$, while the power iteration only scales with $\mathcal{O}(kmn)$ where $k$ is the reduced rank and $k\ll m$.
@@ 4.1 Setup
Datasets. We select four visual classification datasets to investigate the task of balancing personalization, generalization and privacy: Caltech101 (Fei-Fei et al., 2004), OxfordPets (Parkhi et al., 2012), OxfordFlowers (Nilsback & Zisserman, 2008) and Food101 (Bossard et al., 2014). We utilize the pathological data split among $10$ clients. Each client model is trained on its local classes, and evaluated on both its local classes for personalization capability and neighbor classes (classes owned by other clients) for generalization capability. We also evaluate the large-scale dataset CIFAR-100 (Krizhevsky et al., 2009) with the Dirichlet data split among $25$ and $50$ clients. The final test accuracy is obtained by averaging the performance across all clients. Details of our implementation and hyperparameters are provided in the appendix. Our code is available at https://anonymous.4open.science/r/DP-FPL-84F8.
@@ 4.2 Performance Results
Improving personalization in private setting. Table 2 shows the average test accuracy over the last $10$ epochs on local classes. In the non-private scenario (i.e., Noise = None), FedOTP shows the highest utility, however, as we add more noise, DP-FPL comes in the first place with the highest local classes accuracy. Under strict privacy levels ($\epsilon=0.01$), we see a noticeable decrease in test accuracy. Nevertheless, DP-FPL still consistently outperforms other baselines across all datasets, showing the robustness of our method even under strictly private conditions. Table 4 shows the overall accuracy utilizing a Dirichlet data split and ResNet50 as the backbone model. As shown in Table 4, DP-FPL demonstrates superior performance compared to baseline methods across all datasets. This shows the effectiveness and applicability of our approach in various complex settings, including different data distributions, models, and number of clients.
@@ 4.3 Ablation Study
Effect of privacy noise $\epsilon$. We study how different privacy level affects the model performance on both local and neighbor classes in Figure 2 (b) and (c). As expected, the local classes accuracy gradually decreases when noise level increases. However, we see some unexpected improvement in neighbor utility when we increase DP noise from $0.4$ to $0.1$. As explained in Section 4.2, certain privacy noise range acts like regularization error that prevents overfitting, resulting in better generalization for higher noise level.
@@ 5 Conclusion
In this paper, we presented a novel approach to address the critical challenges of personalization, generalization, and privacy in FPL for multimodal LLMs. Our proposed framework leverages low-rank decomposition to balance the tradeoffs between these competing objectives. By decomposing the local prompts into low-rank components iteratively while incorporating a residual term, our method effectively preserves both generalization and personalization. Moreover, we introduced a privacy-preserving mechanism that applies both global and local DP to safeguard sensitive client data. Unlike conventional methods, we selectively applied DP noise to the low-rank components, allowing us to maintain privacy without significantly degrading model performance. The critical role of the residual term in mitigating the effects of DP noise was also demonstrated, highlighting its importance for maintaining model expressiveness and personalization.
@@ A.2 Implementation Details
For the prompt learner, the length of prompt vectors is $b=16$ with a dimension of $d=512$, and the token position is “end” with “random” initialization. For the factorization process, we experiment with four different decomposition rank $1,2,4,8$. We consider three different DP noises with privacy level from low to high: $\epsilon\in\{0.4,0.2,0.1\}$. The clipping threshold is chosen to be $C_{th}=10$ for both GDP and LDP applications.
@@
We run our experiment on a computer cluster, each node has a 6x NVIDIA Tesla V100 GPUs with 32 GiB of memory and 512 GiB RAM and 2x IBM Power 9 processors. The final result is the mean accuracy over the last $10$ training epochs, averaged across all clients.

### v3

@@ Abstract
Multimodal Large Language Models (LLMs) are pivotal in revolutionizing customer support and operations by integrating multiple modalities such as text, images, and audio. Federated Prompt Learning (FPL) is a recently proposed approach that combines pre-trained multimodal LLMs such as vision-language models with federated learning to create personalized, privacy-preserving AI systems. However, balancing the competing goals of personalization, generalization, and privacy remains a significant challenge. Over-personalization can lead to overfitting, reducing generalizability, while stringent privacy measures, such as differential privacy, can hinder both personalization and generalization. In this paper, we propose a Differentially Private Federated Prompt Learning (DP-FPL) approach to tackle this challenge by leveraging a low-rank factorization scheme to capture generalization while maintaining a residual term that preserves expressiveness for personalization. To ensure privacy, we introduce a novel method where we apply local differential privacy to the two low-rank components of the local prompt, and global differential privacy to the global prompt. Our approach mitigates the impact of privacy noise on the model performance while balancing the tradeoff between personalization and generalization. Extensive experiments demonstrate the effectiveness of our approach over other benchmarks.
@@ 1 Introduction
To tackle the above question, we proposed a Differentially Private Federated Prompt Learning (DP-FPL) approach that leverages low-rank factorization and DP as part of the prompt learning process. In our framework, each client simultaneously learns a global prompt and a local prompt. The global prompt is shared in a FL manner for generalized knowledge transfer, while the local prompt is retained at each client site for personalization. Our contributions are threefold.
  • We propose a privacy-preserving personalized federated prompt learning approach with Differential Privacy for multimodal LLMs. We factorize the local prompt into two lower rank components with an additional residual term. The factorized low-rank components allow the model to capture broader patterns that are beneficial across different data distributions, aiding the generalization capability of each client. The residual term is crucial for retaining the expressiveness lost during the factorization process, thereby preserving the client-specific learning and improving personalization.
@@ 2 Related Work
Personalized Federated Learning. There are several existing approaches that aim to learn personalized models for clients in FL settings, including clustering (Ghosh et al., 2020; Berlo et al., 2020; Shahid et al., 2021), regularization (Shoham et al., 2019; Dinh et al., 2020; Li et al., 2020) and knowledge distillation (Li & Wang, 2019; He et al., 2020; Fang & Ye, 2022). Personalized FL is most commonly approached as a multi-task learning problem that simultaneously learns two models for each client: a global model for generalized knowledge and a local model for personalized data. Existing methods accomplish this by decoupling the model parameters or layers into global and local learning components (Arivazhagan et al., 2019; Deng et al., 2020; Zhang et al., 2020; Collins et al., 2021; Jeong & Hwang, 2022; Zhang et al., 2023). In the existing literature on personalized FL, private multi-task learning approaches aim to protect training data by retaining personalized parameters and sharing differentially private generalized parameters. Examples include Jain et al. (2021), Hu et al. (2021), Bietti et al. (2022), Yang et al. (2023b) and Xu et al. (2024). However, these methods are designed for full model training and cannot be directly applied to prompt tuning due to the difference in the parameter space. Sun et al. (2024) incorporates Low-Rank Adaptation with DP in a standard FL setting, but they do not consider personalization and prompt learning, making their method not applicable to our setting.
@@ 3 Proposed Method
We introduce our proposed method, Differentially Private Federated Prompt Learning (DP-FPL), shown in Figure 1. Our approach leverages low-rank factorization with an additional residual term to balance personalization and generalization in a differentially private FPL system.
@@ 3.1 Preliminaries on Personalized Federated Prompt Learning
A simple example of the personalized prompt is illustrated in Figure 1, in which each client has a collection of images captured from different angles, reflecting the heterogeneity in their local data. Consequently, client $1$ might use the prompt ”an upside-down photo of [class]”, while client $i$ could have the prompt ”a upright photo of [class]”, and client $N$ might use ”a rotated photo of [class]”. In this instance, the template ”a photo of [class]” is the generalized global prompt shared across clients, and the terms ”upside-down”, ”upright” and ”rotated” represent the personalized characteristics of each client’s prompt. These variations in prompts allow the client model to better adapt to the specific characteristics of their data, improving personalization.
@@ 3.2 Balancing Personalization and Generalization
Prior research has demonstrated that fine tuning pre-trained LLMs with lower dimension reparameterization promotes generalization capability across various tasks (Aghajanyan et al., 2020; Cui et al., 2024). Additionally, lower intrinsic dimension improves the model utility under the effect of DP noise (Yu et al., 2021; Xu et al., 2024). Therefore, we utilize low-rank factorization as part of our FPL framework to balance the personalization and generalization learning under the influence of DP noise. However, low-rank training such as Low-Rank Adaptation has difficulty matching the performance of full-rank training in many difficult tasks (Liu et al., 2024; Biderman et al., 2024; Ivison et al., 2023; Zhuo et al., 2024). This is because low-rank factorization methods restrict the parameter space, removing some of the information of the full-rank space (Konečnỳ, 2016). Recent work Cui et al. (2024) also used the low-rank approximation in non-private FPL systems, however, they factorize the local prompt only once at the beginning and train iteratively with the low-rank components. This approach can reduce the overall expressive power over the training phase. As we will show in Section 4, the loss of expressiveness can amplify the adverse effects of the DP noise, further diminishing the model’s performance.
To overcome these issues of low-rank training, we perform the factorization process in every training round rather than only at the beginning like Cui et al. (2024), and we incorporate a residual term to compensate for the lost expressiveness. We analyze in detail the benefit of the residual term in Section 3.3. The overall personalized prompt of client $i$ can then be expressed as $p_{i}=p_{G,i}+u_{i}v_{i}+r_{i}$, where $u_{i}$ and $v_{i}$ are the factorized low-rank components and $r_{i}$ is the additional residual term.
We utilize a version of the Reparametrized Gradient Perturbation (RGP) method introduced in Yu et al. (2021) as our low-rank factorization scheme. We perform the factorization process for each client local prompt $p_{L,i}$ in every training iteration to get the temporary low-rank prompt components $u_{i}$ and $v_{i}$. We also compute a residual term $r_{i}$ which is the remainder of the factorization process, i.e. $r_{i}=p_{L,i}-u_{i}v_{i}$. As shown by Yu et al. (2021), each client can reconstruct the gradient of $p_{L,i}$ using the gradient of the low-rank terms in back propagation as follows:
@@ 3.3 Preserving Privacy
The two main building blocks in our approach, i.e., low-rank factorization and DP, naturally introduce error to the training process. We conjecture that this error acts as a regularization term that prevents clients from overfitting to local data, reducing personalization and improving generalization. However, under strictly private conditions (lower rank and higher DP noise), the accumulated error may become too large and potentially destroy the personalization capability. In this case, the added residual term compensates for the regularization-like error and helps improve local learning, balancing personalization and generalization. We demonstrate the benefit of the residual term in the ablation study in Section 4.3, supporting our hypothesis. Further theoretical analysis of the residual term can be a potential future work direction.
@@ 3.4 Algorithm
Low-rank factorization via traditional SVD method requires significant runtime. Instead, we use the power method with one iteration, significantly reducing the computational cost. Given the full-rank matrix of size $m\times n$ (assuming $m\leq n$), the computational cost of SVD scales with $\mathcal{O}(m^{2}n)$, while the power iteration only scales with $\mathcal{O}(kmn)$ where $k$ is the reduced rank and $k\ll m$.
@@ 4.1 Setup
Datasets. We select four visual classification datasets to investigate the task of balancing personalization, generalization and privacy: Caltech101 (Fei-Fei et al., 2004), OxfordPets (Parkhi et al., 2012), OxfordFlowers (Nilsback & Zisserman, 2008) and Food101 (Bossard et al., 2014). We utilize the pathological data split among $10$ clients. Each client model is trained on its local classes, and evaluated on both its local classes for personalization capability and neighbor classes (classes owned by other clients) for generalization capability. We also evaluate the large-scale dataset CIFAR-100 (Krizhevsky et al., 2009) with the Dirichlet data split among $25$ and $50$ clients. The final test accuracy is obtained by averaging the performance across all clients. Details of our implementation and hyperparameters are provided in the appendix, and the source code is publicly available  [^1] https://github.com/linhhtran/DP-FPL.
@@ 4.2 Performance Results
Improving personalization in private setting. Table 2 shows the average test accuracy over the last $10$ epochs on local classes. In the non-private scenario (i.e., Noise = None), FedOTP shows the highest utility, however, as we add more noise, DP-FPL has the highest local classes accuracy. Under strict privacy levels ($\epsilon=0.01$), we see a noticeable decrease in test accuracy. Nevertheless, DP-FPL still consistently outperforms other baselines across all datasets, showing the robustness of our method even under strictly private conditions. Table 4 shows the overall accuracy utilizing a Dirichlet data split and ResNet50 as the backbone model. As shown in Table 4, DP-FPL demonstrates superior performance compared to baseline methods across all datasets. This shows the effectiveness and applicability of our approach in various complex settings, including different data distributions, models, and number of clients.
@@
Additional results. We include additional experiment on the performance of MIA against DP-FPL in the appendix (see Section A.3). The results show that the attack success rate is relatively low when $\epsilon=0.1$ for all datasets. In addition, $\epsilon=0.1$ causes less than $10\%$ reduction in the target model accuracy for both local and neighbor classes as shown in Tables 2 and 3.
@@ 4.3 Ablation Study
Effect of privacy noise $\epsilon$. We study how different privacy level affects the model performance on both local and neighbor classes in Figure 2 (b) and (c). As expected, the local classes accuracy gradually decreases when noise level increases. However, we see some unexpected improvement in neighbor utility when we increase DP parameter $\epsilon$ from $0.4$ to $0.1$. As explained in Section 4.2, certain privacy noise range acts like regularization error that prevents overfitting, resulting in better generalization for higher noise level.
@@ 5 Conclusion
In this paper, we presented a novel approach to address the critical challenges of personalization, generalization, and privacy in FPL for multimodal LLMs. Our proposed framework leverages low-rank factorization to balance the tradeoffs between these competing objectives. By factorizing the local prompts into low-rank components iteratively while incorporating a residual term, our method effectively preserves both generalization and personalization. Moreover, we introduced a privacy-preserving mechanism that applies both global and local DP to safeguard sensitive client data. Unlike conventional methods, we selectively applied DP noise to the low-rank components, allowing us to maintain privacy without significantly degrading model performance. The critical role of the residual term in mitigating the effects of DP noise was also demonstrated, highlighting its importance for maintaining model expressiveness and personalization.
@@ A.2 Implementation Details
For the prompt learner, the length of prompt vectors is $b=16$ with a dimension of $d=512$, and the token position is “end” with “random” initialization. For the factorization process, we experiment with four different factorization rank $1,2,4,8$. We consider three different DP noises with privacy level from low to high: $\epsilon\in\{0.4,0.2,0.1\}$. The clipping threshold is chosen to be $C_{th}=10$ for both GDP and LDP applications.
@@
We run our experiment on a computer cluster, each node has a 6x NVIDIA Tesla V100 GPUs with 32 GiB of memory and 512 GiB RAM and 2x IBM Power 9 processors. The final result is the mean accuracy across all clients, averaged over $5$ runs with different seeds.